OpenStudy (anonymous):

2.) P is 9 more than half of q.

OpenStudy (kuoministers):

you just need to formulate the equation

OpenStudy (kuoministers):

\[\frac{ 1 }{ q } + 9\] would you agree that this is correct?

OpenStudy (kuoministers):

or this?

OpenStudy (kuoministers):

\[\frac{ q }{ 2 } + 9\]

OpenStudy (kuoministers):

\[p = \frac{ q }{ 2 } + 9\]